### v1

# A Tidy Data Structure and Visualisations for Multiple Variable Correlations and Other Pairwise Scores

Abstract
We provide a pipeline for calculating, managing and visualising correlations and other pairwise scores for numerical and categorical data. We present a uniform interface for calculating a plethora of pairwise scores and a new tidy data structure for managing the results. We also provide new visualisations which simultaneously show multiple and/or grouped pairwise scores. The visualisations are far richer than a traditional heatmap of correlation scores, as they help identify relationships with categorical variables, numeric variable pairs with non-linear associations or those which exhibit Simpson’s paradox. These methods are available in our R package bullseye.
1   Introduction
Correlation visualisation, frequently a heatmap matrix, is a common first step in investigating relationships between variables, especially in high-dimensional datasets. Typically, Pearson’s correlation for numeric variables is used, but in principle the same visualisations could be used for measures of association which assess non-linear patterns, and for summarising relationships for non-numeric variables. Heatmap visualisations show a single bivariate summary for each pair of variables, but richer visualisations could be designed to show multiple summaries, capturing a variety of bivariate patterns. Such visualisations would also be useful for investigating group-wise correlations and association measures.
In this paper we present a new package bullseye (Chinwan and Hurley 2024) which offers a uniform interface for calculating various scores assessing the relationship between variable pairs, both numerical and categorical. We construct a tidy data structure for managing multiple and/or group-wise bivariate scores. We design a new visualisation method which is a variant of the correlation heatmap matrix for showing multiple pairwise or grouped scores. The variant which compares overall to grouped scores is reminiscent of a bullseye, hence the package name. We further design a new linear display, which is useful for higher-dimensional datasets, when interest lies in displaying just interesting pairs of variables. Both displays are optionally interactive.
More general pairwise variable scores known as scagnostics also fit into the framework offered by bullseye. These numeric variable scores were first introduced by Tukey and Tukey (1985) and further developed by Wilkinson, Anand, and Grossman (2005) and seek to identify patterns such as clumpiness and the presence of outliers, for the purpose of identifying scatterplots worthy of detailed exploration.
The outline of the paper is as follows: In Section 2 we present background, first discussing R packages offering visualisations of correlation summaries, and then move to R packages providing calculation of pairwise variable scores beyond the familiar Pearson, Spearman or Kendall’s correlation. In Section 3 we introduce bullseye, describing the new tidy data structure for managing pairwise scores, and the range of functions offered for calculating them. In Section 4 we present the new visualisations offered in bullseye and give some examples. Section 5 describes how bullseye can be used in co-operation with other R related packages. Finally we conclude with a discussion.
2   Related R packages
In this section, we provide a brief review of existing R packages used for correlation displays. This establishes the gap that the current work aims to fill. We then provide an overview of methods for calculating pairwise association for numerical and factor variables, specifically with reference to their R package implementations
Displaying correlation
There are many available R packages providing correlation displays, the most notable are summarized in Table 1. For each package we list the display type offered, whether categorical variables are supported and whether grouped correlations are on offer. Most of these use a matrix layout of glyphs or filled squares (i.e., a heatmap) and cater for numerical variables only.
Table 1: List of R packages for correlation display, with the display type, whether categorical variables are supported, whether grouped correlations are provided, and whether a dataframe is used to represent correlations.
package Display type Categorical Groups Dataframe
ellipse ellipse matrix
corrplot heatmap/glyph matrix
corrgram heatmap/glyph matrix
ggcorrplot heatmap/glyph matrix
corrr glyph matrix/network
linkspotter network
corrgrapher network
correlation heatmap
correlationfunnel linear
D. J. Murdoch and Chow (1996) proposed a display for correlation matrices which uses a matrix layout of ellipses where the parameters of the ellipses are scaled to the correlation values. An implementation is provided in package ellipse (D. Murdoch and Chow 2023). Friendly (2002) expanded on this idea by rendering correlation values as shaded squares, bars, ellipses, or circular pac-man symbols, and these methods are implemented in packages corrplot (Wei and Simko 2021) and corrgram (Wright 2021). The ggcorrplot (Kassambara 2023) and corrr (Kuhn, Jackson, and Cimentada 2024) packages also provided heatmap-type displays, but based on ggplot2 (Wickham 2016) rather than base R.
A number of packages support network displays of correlation values. In the package corrr network displays use line thickness to encode correlation magnitude, with a filtering option to discard low-correlation edges. linkspotter (Samba 2020) offers a variety of association measures in addition to correlation, where the measure used depends on whether the variables are both numerical, categorical or mixed. The results are visualised in a network plot, which may be packaged into an interactive shiny application. The package corrgrapher (Morgen and Biecek 2020) also uses a network plot, and handles mixed type variables by using association measures obtained as transformations of $p$-values obtained from Pearson’s correlation test in the case of two numeric variables, Kruskal’s test for numerical and factor variables, and a chi-squared test for two categorical variables.
The R package correlationfunnel Dancho (2020) offers a novel linear display which assists in feature selection in a regression setting with many predictor variables. All numeric variables including the response are binned. All (now categorical) variables in the resulting dataset are one-hot encoded and Pearson’s correlation calculated with the response categories. The correlations are visualised in a dot-plot display, where predictors are ordered by maximum correlation magnitude, giving a funnel-like shape.
The majority of the listed packages display the result of the standard cor calculation, so the input data structure for the visualisations is a $p\times p$ symmetric matrix of numerical values, where $p$ is the number of variables. corrr uses a $p\times p$ dataframe instead, or alternatively a long format dataframe where each pair of variables form a row, and provides some convenience functions for manipulating these structures. linkspotter similarly uses a long dataframe where pairs of variables form a row, with additional columns for possible multiple correlations. The package correlation (Makowski et al. 2022) also uses a long dataframe, and for grouped correlations each row is associated with a variable pair and group level. It also offers heatmap and network-style visualisations, though not, as far as we can tell, for grouped correlations.
Calculating association measures
Many alternative measures of association have been devised, beyond the standard Pearson, Spearman or Kendall’s correlation. The purpose of these measures is to capture non-linear associations, associations between nominal and ordinal variables, as well as associations between variables of mixed type. Implementations of these measures are scattered across many R packages.
For nominal variable pairs, many classical measures of association are implemented in the R package DescTools (Signorell 2024). One of these is Pearson’s contingency coefficient which uses the ${\chi}^{2}$ value (scaled to the unit interval) from Pearson’s ${\chi}^{2}$ test for independence. Another is the uncertainty coefficient (Theil 1970) which measures the proportion of uncertainty in one variable which is explained by the other. A symmetric version of this coefficient is obtained by taking the mean of the uncertainty coefficients treating each variable as an independent variable once.
DescTools also offers implementations of many association measures for ordinal pairs of variables, including Kendall’s $\tau$ and $\tau_{b}$ and Goodman Kruskal’s gamma, all of which are based on the on the number of concordances and discordances in paired observations (see for example Agresti (2010) for a description). polycor (Fox 2022) implements polychoric correlation (Olsson 1979), which summarizes the association between two ordinal variables by assuming two normally distributed latent variables. It also provides polyserial correlation, a measure of association between an ordinal, numeric pair of variables.
For numeric pairs of variables, the recently proposed measures distance correlation (Székely, Rizzo, and Bakirov 2007) and maximal information coefficient (MIC) (Reshef et al. 2011) are measures of dependence which aim to capture non-linear as well as linear patterns. Distance correlation is calculated as the correlation between two sets of distances, and is available from the R package energy (Rizzo and Szekely 2022). The MIC coefficient is calculated as the mutual information between discretised versions of the two numeric variables, where the binning is chosen to maximise the mutual information. Simon and Tibshirani (2014) simulated pairs of variables with different relationships at varying levels of noise and showed that distance correlation has more statistical power than MIC. In a later study, Kinney and Atwal (2014) proposed additional mutual information-based measures with improved performance. The MIC and other maximal information based non-parametric exploration statistics have been implemented in the R package minerva (Albanese et al. 2012). Another package linkspotter provides a mutual information measure for pairs of variables which may both be numeric, factors or mixed, which is similar to the minerva MIC measure for numeric pairs.
The alternating conditional expectations (ACE) algorithm (Breiman and Friedman 1985) as implemented in acepack (Spector et al. 2016) estimates optimal transformations between response and predictor variables for regression analysis and calculates $R^{2}$ using this optimal transformation. For a numeric pair of variables, this provides another way of summarizing their non-linear association. For categorical variables, ACE transformations calculates the optimal scoring of levels (also known as canonical analysis), and the resulting $R^{2}$ provides an alternative measure of association.
The package scagnostics (Wilkinson and Anand 2022) implements the methods in Wilkinson, Anand, and Grossman (2005) evaluating scatterplot “interestingness”, and provides nine scores called outlying, skewed, clumpy, sparse, striated, convex, skinny, stringy and monotonic.
3   Bullseye calculations
In this section and the following Section 4 we describe the main features of our new package bullseye. Here we present a tidy data structure for managing pairwise scores, which offers far more flexibility than the standard correlation matrix. Then we describe the functions included for calculating individual, multiple and grouped pairwise scores. In the next section we discuss the new visualisations offered and give some examples.
For the examples in this section we use the penguins dataset from package palmerpenguins (Horst, Hill, and Gorman 2020), which lists measurements for penguin species from three islands in the Palmer Archipelago. The dataset has five numerical variables, and three factor variables species, island and sex.
Design of tidy data structure
Conventionally, correlation measures are represented as a matrix, following the mathematical definition of a correlation matrix. This matrix representation does not follow tidy data principles and does not extend naturally to the situation of multiple or grouped pairwise measures. For these reasons, some existing packages (for example corrr, linkspotter and correlation) provide tidy dataframe representations.
For our purposes, we design a pairwise structure that contains as columns (i) the two variables, (ii) the score name (iii) the group (iv) the score value and (v) the types of the variables. Starting from a correlation matrix of numeric variables in the penguins dataset, constructed below as peng_cor, the pairwise function constructs the equivalent pairwise structure:
peng_cor <- select(penguins, where(is.numeric)) |>
  cor(use= "pairwise.complete.obs")
pairwise(peng_cor, score="pearson", pair_type = "nn")
#> # A tibble: 10 x 6
#>   x         y         score   group  value pair_type
#>   <chr>     <chr>     <chr>   <chr>  <dbl> <chr>
#> 1 bill_dep  bill_len  pearson all   -0.235 nn
#> 2 bill_len  flip_len  pearson all    0.656 nn
#> 3 bill_dep  flip_len  pearson all   -0.584 nn
#> 4 body_mass flip_len  pearson all    0.871 nn
#> 5 bill_len  body_mass pearson all    0.595 nn
#> 6 bill_dep  body_mass pearson all   -0.472 nn
#> # i 4 more rows
As bullseye is designed for symmetric measures, each variable pair appears once only. By construction the x and y variables are in alphabetical order, that is, x < y. Note there is no row where x = y. Here the correlations are not grouped so group="all". In the pair_type column, we use nn for two numeric variables, ff for two factors and fn for a factor-numeric pair. The above calculation can be done in one step with the bullseye function pair_cor. Notice that this automatically excludes non-numeric variables and handles missing values using pairwise complete observations. In the example below, the optional argument method = "spearman" is specified so Spearman correlation is calculated, rather than the default Pearson.
pair_cor(penguins, method="spearman")
#> # A tibble: 10 x 6
#>   x         y         score    group  value pair_type
#>   <chr>     <chr>     <chr>    <chr>  <dbl> <chr>
#> 1 bill_dep  bill_len  spearman all   -0.222 nn
#> 2 bill_len  flip_len  spearman all    0.673 nn
#> 3 bill_dep  flip_len  spearman all   -0.523 nn
#> 4 body_mass flip_len  spearman all    0.840 nn
#> 5 bill_len  body_mass spearman all    0.584 nn
#> 6 bill_dep  body_mass spearman all   -0.432 nn
#> # i 4 more rows
Calculating pairwise scores
Table 2 has a list of functions provided by bullseye to calculate pairwise variable scores. Each of these functions takes as input a data.frame containing a dataset and handle.na which defaults to TRUE (and uses pairwise complete observations). Each excludes variables which are not of the appropriate type for the score to be calculated. Other arguments may optionally be passed to the underlying calculation function, specified in the “From” column of Table 2. Each of these pair_xx functions produces an output which is of class pairwise, with the score and pair_type columns filled in as appropriate. The function pair_mine will optionally provide multiple measures for each variable pair, while pair_scagnostics defaults to providing all nine scagnostics from scagnostics.
Table 2: Bullseye functions for calculating pairwise corrrelations, other association measures and scagnostics. Tick marks in the columns nn, ff, fn indicate whether the function gives a score for numeric, factor and factor-numeric pairs. Ticks in the Ordinal column indicates method assumes ordinal factors.
Function Description nn ff fn From Range Ordinal
pair_cor Pearson/Spearman/Kendall cor [-1,1]
pair_dcor Distance correlation energy::dcor2d [0,1]
pair_mine MIC and other measures minerva::mine [0,1]
pair_ace Ace correlation acepack::ace [0,1]
pair_cancor Canonical correlation cancor [0,1]
pair_nmi MIC linkspotter::maxNMI [0,1]
pair_polychor Polychoric correlation polycor::polychor [-1,1]
pair_polyserial Polyserial correlation polycor::polyserial [-1,1]
pair_tauA Kendall’s tau A DescTools::KendalTauA [-1,1]
pair_tauB Kendall’s tau B DescTools::KendalTauB [-1,1]
pair_tauC Stuart-Kendall Tau-C DescTools::StuartTauC [-1,1]
pair_tauW Kendall’s W DescTools::KendalW [-1,1]
pair_gkGamma Goodman-Kruskall Gamma DescTools::GoodmanKruskalGamma [-1,1]
pair_gkTau Goodman-Kruskall tau DescTools::GoodmanKruskalTau [0,1]
pair_uncertainty Uncertainty coefficient DescTools::UncertCoef [0,1]
pair_chi Pearson’s contingency coefficient DescTools::ContCoef [0,1]
pair_scagnostics Scagnostics scagnostics::scagnostics [0,1]
We previously saw an example of pair_cor in action. Here is pair_mine, where two methods are specified. The default choice here is just method="MIC".
pair_mine(penguins, method=c("MIC", "TIC"))
#> # A tibble: 20 x 6
#>   x        y         score group value pair_type
#>   <chr>    <chr>     <chr> <chr> <dbl> <chr>
#> 1 bill_dep bill_len  MIC   all   0.313 nn
#> 2 bill_dep bill_len  TIC   all   0.236 nn
#> 3 bill_dep body_mass MIC   all   0.518 nn
#> 4 bill_dep body_mass TIC   all   0.435 nn
#> 5 bill_dep flip_len  MIC   all   0.660 nn
#> 6 bill_dep flip_len  TIC   all   0.510 nn
#> # i 14 more rows
Additional bullseye functions provide multiple pairwise variable scores and/or grouped scores, and these are listed in Table 3.
Table 3: Bullseye functions for combining multiple pairwise scores, and for handling grouping variables.
Function Description
pairwise_multi Combines outputs of pair_xx functions.
pairwise_by Combines outputs of a pair_ xx function for every level of a grouping variable.
pairwise_scores Combines outputs of pair_ xx functions, where different pair_xx functions may be specified for nn, fn, and ff pairs. A grouping variable may optionally be specified.
Dataframes of type pairwise may of course be bound together rowwise. For convenience, we provide another function pairwise_multi which calculates multiple association measures for every variable pair in a dataset. The example below calculates distance correlation and MIC for penguin variables.
dcor_nmi <- pairwise_multi(penguins, c("pair_dcor", "pair_nmi"))
For each of the pairwise calculation functions, they can be wrapped using pairwise_by to build a score calculation for each level of a grouping variable. (Of course, grouped scores could be calculated using dplyr (Wickham et al. 2023) machinery, but it is a bit more work.) In the code below, correlations are calculated for each penguin species, in addition to the overall or ungrouped correlation. Supplying argument ungrouped=FALSE suppresses calculation of the ungrouped scores.
pairwise_by(penguins, by="species", pair_cor)
#> # A tibble: 40 x 6
#>   x        y         score   group      value pair_type
#>   <chr>    <chr>     <chr>   <fct>      <dbl> <chr>
#> 1 bill_dep bill_len  pearson Adelie     0.391 nn
#> 2 bill_dep bill_len  pearson Chinstrap  0.654 nn
#> 3 bill_dep bill_len  pearson Gentoo     0.643 nn
#> 4 bill_dep bill_len  pearson all       -0.235 nn
#> 5 bill_dep body_mass pearson Adelie     0.576 nn
#> 6 bill_dep body_mass pearson Chinstrap  0.604 nn
#> # i 34 more rows
Finally, the function pairwise_scores calculates a different score depending on variable types:
pairwise_scores(penguins) |>
  distinct(score, pair_type)
#> # A tibble: 3 x 2
#>   score   pair_type
#>   <chr>   <chr>
#> 1 cancor  ff
#> 2 cancor  fn
#> 3 pearson nn
The default uses Pearson’s correlation for numeric pairs, and canonical correlation for factor-numeric or factor-factor pairs. In addition, polychoric correlation is used for two ordered factors, but there are no ordered factors in this data. Alternative scores may be specified using the control argument to pairwise_scores. pairwise_scores also has a by argument, and in this way provides pairwise scores for the levels of a grouping variable. Once again, use ungrouped=FALSE to suppress calculation of the ungrouped scores.
pairwise_scores(penguins, by="species")
#> # A tibble: 84 x 6
#>   x         y      score  group   value pair_type
#>   <chr>     <chr>  <chr>  <fct>   <dbl> <chr>
#> 1 island    sex    cancor Adelie 0.0164 ff
#> 2 bill_len  island cancor Adelie 0.0838 fn
#> 3 bill_dep  island cancor Adelie 0.0629 fn
#> 4 flip_len  island cancor Adelie 0.148  fn
#> 5 body_mass island cancor Adelie 0.0208 fn
#> 6 bill_len  sex    cancor Adelie 0.590  fn
#> # i 78 more rows
The package has a tibble called pair_methods which lists the information provided in Table 2. This makes it easy to search for methods suitable for particular variable types. Below we identify the methods suitable for numeric, factors and combined numeric-factor pairs.
filter(pair_methods, nn&ff&fn)
#> # A tibble: 3 x 7
#>   name        nn    ff    fn    from                range ordinal
#>   <chr>       <lgl> <lgl> <lgl> <chr>               <chr> <lgl>
#> 1 pair_ace    TRUE  TRUE  TRUE  acepack::ace        [0,1] FALSE
#> 2 pair_cancor TRUE  TRUE  TRUE  cancor              [0,1] FALSE
#> 3 pair_nmi    TRUE  TRUE  TRUE  linkspotter::maxNMI [0,1] FALSE
4   Bullseye visualisations
In this section we describe the visualisation functions plot_pairwise and plot_pairwise_linear offered by bullseye. Then we present some examples.
Visualisation functions
The first visualisation function plot_pairwise is the default plot method for objects of class pairwise. It provides a traditional matrix layout, generalising the usual heatmap. For each (x,y) pair a circular glyph is drawn, divided into equal-sized wedges if the (x,y) pair appears multiple times. Each wedge has a fill level according to the associated value. To distinguish rows representing the complete dataset from those representing subsets, the circular glyph is divided into an inner circle (the bullseye) and an outer doughnut. Wedges in the inner circle represent the complete dataset (identified as rows with group = "all"), and segments of the doughnut represent dataset subsets. In an earlier variant of this display (see Chinwan (2024) Figure 3.4) , we used lollipops to show the scores in each matrix cell, but we find the bullseye version more impactful, though admittedly this assessment was not the subject of an experiment.
Seriation is used to order variable pairs to help draw attention to high-score pairs. In the case of a single score for each (x,y) pair, the usual way of ordering variables employs hierarchical clustering to place high-score (typically correlation) pairs nearby. This has the effect of placing high-score pairs of variables near the diagonal of the matrix layout. In plot_pairwise we use a variant of this scheme, which additionally aims to give high scoring pairs extra prominence by placing them in the top-left corner, using the Earle and Hurley (2015) lazy path length algorithm implemented in package DendSer (dendser?). For multiple scores, for each (x,y) pair we use the maximum of the absolute values, or alternatively the maximum difference.
The second visualisation function plot_pairwise_linear is also available using the plot method for objects of class pairwise, but with type="linear". Each row of the display shows all the values for a given (x,y) pair, using either a tile or point geom. This display is useful for datasets with lots of variables, where the pairwise structure can optionally be filtered to show interesting (x,y) pairs only, or in the case of a regression problem, pairs where a specified response variable is present. Once again, seriation is used to order variable pairs to help draw attention to high-score pairs. Seriation for this linear display is more straightforward than that for the matrix display. In plot_pairwise_linear variable pairs are ordered by their maximum absolute value, or alternatively the maximum difference.
Both visualisations functions produce ggplot2 (ggplot2) objects (Wickham 2016). Alternatively, they have an interactive option, in which case package ggiraph (Gohel and Skintzos 2024) is used and tooltips provide more information on geom objects.
Visualisation examples
We start by displaying pairwise association measures for all variables in the penguins data.
plot(pairwise_scores(penguins))
Figure 1: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs. Notice the categorical variable species is strongly associated with the penguin dimension variables, and with the categorical variable island.
The above code calculates pairwise_scores for the penguins data, Figure 1 shows the resulting display. As the variables species, sex and island are factors, scores involving any of these variables is a canonical correlation, while Pearson’s correlation is shown for numerical pairs. As the species variable is strongly associated with the measurement variables, it is clear that the size varies across the penguin species. The species variable is also strongly associated with island, telling us that the species distribution varies across the islands. Not surprisingly, year has little association with the other variables.
Next, we investigate the associations for all variables grouped by species. The code below uses pairwise_scores to calculate pairwise associations grouped by species. We use var_order="seriate_max_diff" to bring variable pairs with the biggest score difference to the top-left corner. Figure 2 shows the result.
sc_species<- pairwise_scores(penguins, by="species")
plot(sc_species, var_order="seriate_max_diff")
Figure 2: Pairwise correlations for numerical variables and canonical correlation for factors and mixed factor-numeric pairs grouped by species. Ungrouped measures are in the centre bullseye, outer wedges represent Adelie, Chinstrap, Gentoo in anti-clockwise order. The grey colour represents NA values. Tooltips show the groups and the score values in the html version. A blue inner circle surrounded by red wedges indicates that direction of correlation is reversed within groups for some variable pairs.
The inner circle of Figure 2 shows the ungrouped score, and the outer doughnut shows the scores for each of the species. The html version of this plot is interactive, which is useful to obtain the species represented in each segment and the numerical value of the score. Generally, it is of interest if there is a high difference among the grouped and ungrouped scores. Here the variable bill_dep has a negative overall correlation with each of flip_len, body_mass, bill_len but positive correlation within each of the species groups. The same pattern (known as Simpson’s paradox) is present for the pair bill_len and bill_dep. There are also a number of wedges coloured in grey involving the variable island, representing score values of NA. This is because two of the species (Gentoo and Chinstrap) are located on one island apiece only. For these species, island is constant and so the association value cannot be calculated and is therefore recorded as NA.
To demonstrate the linear display available in bullseye, we switch to a dataset with more variables. We use the acs12 dataset from the package openintro (Çetinkaya-Rundel et al. 2024) which contains results from the US Census American Community Survey in 2012. The dataset has 13 variables, only four of which (income, hrs_work, age, time_to_work) are numeric, the rest are factors. The code segment below calculates ace, cancor and nmi scores. (These three measures were chosen because they are applicable to both numerical and factor variables.)
sc_multi <- pairwise_multi(acs12, scores=c("pair_cancor", "pair_ace", "pair_nmi"))
A display similar to Figure 2 without the center segments could be constructed with the code
plot(sc_multi, interactive=TRUE)
However, rather than displaying all (${{13}\choose{2}}=78$ variable pairs we chose instead to filter out the interesting variable pairs from sc_multi. In this example, we show variable pairs whose maximum score across the three measures exceeds 0.25. The filtered pairwise structure no longer has all (x,y) pairs, so we use the linear display as shown in Figure 3 instead of the matrix display.
sc_multi |>
  mutate(valmax = max(abs(value)), .by=c(x,y))|>
  filter(valmax > .25) |>
  plot(type="linear")
Figure 3: Canonical correlation, ace and nmi measures for the acs12 data. Only variable pairs with scores above 0.25 are shown. Tooltips show the score values in the html version. The canonical age-income correlation is close to zero but the maximal ace correlation is 0.47, indicating a non-linear association.
In Figure 3 variable pairs are sorted from highest to lowest maximum score. We see that the mutual information measure (nmi) is consistently less than the ace score. By definition the ace score cannot be less than the cancor score, and they are identical for categorical pairs of variables, for example, (citizen, lang) and (lang, race). The ace scores for income with other variables is much higher than the cancor score, indicating non-linear association. For example, an investigation of the (age, income) scatterplot shows that income goes up with age until about age 40 and then drops off.
Next we investigate how the variable associations vary by race. We use Pearson correlation for numerical variables and canonical correlation otherwise. Here we filter out those with moderate correlations or those whose correlation varies across race. See the code segment below. Some interesting patterns emerge, see Figure 4.
group_scores <- pairwise_scores(acs12, by = "race")
rng <- function(vals){
  if (all(is.na(vals))) 0 else max(vals, na.rm=TRUE)- min(vals,na.rm=TRUE)
mutate(group_scores, valrange = rng(value),valmax = max(abs(value)), .by=c(x,y))|>
  filter(valrange > .25 | valmax > .5 ) |>
  plot(type="linear", geom="point", pair_order = "seriate_max_diff")+
   theme(legend.text = element_text(size = rel(.5)),
         legend.title = element_text(size = rel(.5))
Figure 4: Pearson/canonical correlations (for numerical/factors) by race for the acs12 data. Only variable pairs with a maximum correlation of 0.5 and a range of 0.25 or above are shown. Associations are highest for Asians in most cases (green dots).
Asians (represented by a green dot) have much higher association than other groups for many pairs of variables. As the pairs are ordered by maximum difference, the plots corresponding to the first few variable pairs are worth further investigation, if difference in correlations across race is of interest. Plotting these variables (not shown), we find that for Asians, those in employment work many more hours per week than do those not in the labour force and unemployed, whereas for other race groups there are much smaller variations in hours worked across employment status. However, there are just a handful of Asians in the data that that were not employed with non-missing values of hrs_work. Also, Asian men have much longer commutes (time_to_work) than Asian women, but for other races male and female commute times are similar. And, the gender wage gap is much higher for Asians (males earning more) than for other groups.
The linear display presented here will also be useful in the regression context to summarize the association patterns between a response and predictor variables, where relationships between predictors is not the focus.
5   Bullseye and other packages
As discussed in Section 2, there are many other R packages which offer displays for correlation and other association measures. Most of these base their displays on a correlation matrix. Our package bullseye has an as.matrix method for pairwise structures, which uses the first entry for each (x,y) pair, so these displays are easily constructed for pairwise objects. Conversely, we provide an as.pairwise method so that any symmetric matrix can be converted to a pairwise class, so bullseye visualisations are accessible.
The correlation package offers calculation of a variety of correlations beyond those currently available in bullseye, including partial correlations, Bayesian correlations, multilevel correlations, biweight, percentage bend or Sheperd’s Pi correlations. The output data structure is a tidy dataframe with a correlation value and correlation tests for variable pairs for which the correlation method is defined. This is converted to pairwise via a provided as.pairwise method, so bullseye visualisations can be used.
Finally, the pairwise structure is easily manipulated to construct general plots: The example code below and resulting display in Figure 5 shows that pivoting the sc_multi pairwise structure containing scores calculated on the acs12 data, gives a dataset where the relationship between the scores can easily be compared using a standard ggplot2 visualisation. Here we see that for these data there is an increasing trend in the relationship between ace and nmi scores, though aces scores are higher.
sc_multi |>
  pivot_wider(names_from=score, values_from = value) |>
  ggplot(aes(x=nmi, y=ace))+
  geom_point()+
  geom_abline(color="blue")
Figure 5: Relationship between nmi and ace measures calculated on the ‘acs12‘ data, overlaid with the line $y=x$. There is an increasing trend between the two measures, though ace scores are higher.
6   Discussion
In this paper we described a new R package bullseye which fills a gap in providing a standard interface for the calculation of (grouped) pairwise variable scores, a tidy data structure for their manipulation, and a new, optionally interactive, visualisation for the display of multiple pairwise scores. As demonstrated in Section 4 the visualisation is particularly useful for easy identification of the presence of Simpson’s paradox. A second linear display is useful to hone in on a filtered set of interesting pairwise scores.
As explained in Section 3, the pairwise constructor assumes that pairwise variable scores are symmetric, and similarly at present the visualisation functions assume symmetry. It would however be straightforward to extend to asymmetric measures and visualisations, and furthermore to include scores and visualisations for single variables.
Of course, there are a plethora of methods for measuring pairwise patterns which are not included in our package. For example, Tjøstheim, Otneim, and Støve (2022) provides a survey of recent developments in measures of dependence for numerical variables. As discussed in Section 4.3, bullseye visualisations are easily accessible for pairwise scores implemented in R beyond those implemented in our package.
The interactive capabilities of bullseye are currently limited to providing informative tooltips. In future, our package could be used as vehicle for interactively displaying interesting pairwise variable plots. Possibilities in this area have been explored by Grimm (2016), who in the archived package mbgraphic, used correlation displays and scagnostics to identify informative scatterplot matrices and parallel coordinate plots embedded in an interactive Shiny application, and also by Dang and Wilkinson (2014), though their work is not implemented in an R package.
7   Acknowledgement
This publication has emanated from research supported by a grant from Science Foundation Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.